A $530,000 trust fund earns 9.9% compounded semi-annually. It is to makA $530,000 trust fund earns 9.9% compounded semi-annually. It is to make payments at the end of every month in perpetuity. What will be the size of the monthly payments? Enter answer to 2 decimal placese payments at the end of every month in perpetuity. What will be the size of the monthly payments? Enter answer to 2 decimal places
Answer #1

Rate Semi annually =9.9%/2
Effective monthly rate =(1+9.9%/2)^(1/6)-1 =0.80848171%
Size of Monthly Payment =Trust fund *Effective monthly rate =530000*0.80848171%=4284.95